Chapter 18

Conclusions and further reading

We have not been exhaustive in the description of the Caml Light features. We only introduced general concepts in functional programming, and we have insisted on the features used in the prototyping of ASL: a tiny model of Caml Light typing and semantics.

The reference manual [20] provides an exhaustive description of the Caml Light language, its libraries, commands and extensions.

Those who read French are referred to [38], a progressive, but thorough introduction to programming in Caml Light, with many interesting examples, and to [21], the French edition of the Caml Light reference manual.

Description about “Caml Strong” and useful information about programming in Caml can be found in [9] and [37].

An introduction to lambda-calculus and type systems can be found in [17], [12] and [4].

The description of the implementation of call-by-value functional programming languages can be found in [19].

The implementation of lazy functional languages is described in [29] (translated in French as [30]). An introduction to programming in lazy functional languages can be found in [5].

134 CHAPTER 18. CONCLUSIONS AND FURTHER READING

Bibliography

[1] H. Abelson and G.J. Sussman. Structure and Interpretation of Computer Programs. The MIT Press, 1985.

[2] A. Aho, R. Sethi, and J. Ullman. Compilers: Principles, Techniques and Tools. Addison Wesley, 1986.

[3] J. Backus. Can programming be liberated from the Von Neumann style? A functional style and its algebra of programs. In Communications of the ACM, volume 21, pages 133–140, 1978.

[4] H.P. Barendregt. The Lambda-Calculus: Its Syntax and Semantics, volume 103 of Studies in Logic and the Foundations of Mathematics. North-Holland, 1984.

[5] R. Bird and P. Wadler. Introduction to Functional Programming. Series in Computer Science. Prentice-Hall International, 1986.

[6] A. Church. The Calculi of Lambda Conversion. Princeton University Press, 1941.

[7] D. Clément, J. Despeyroux, T. Despeyroux, and G. Kahn. A simple applicative language: Mini-ML. In Proceedings of the ACM International Conference on Lisp and Functional Programming, pages 13–27, 1986.

[8] G. Cousineau, P.-L. Curien, and M. Mauny. The categorical abstract machine. In Functional Programming Languages and Computer Architecture, number 201 in Lecture Notes in Computer Science, pages 50–64. Springer Verlag, 1985.

[9] G. Cousineau and G. Huet. The CAML primer. Technical Report 122, INRIA, 1990.

[10] N. De Bruijn. Lambda-calculus notation with nameless dummies, a tool for automatic formula manipulation. Indag. Math., 1962.

[11] M. Gordon, R. Milner, L. Morris, M. Newey, and C. Wadworth. A metalanguage for interactive proofs in LCF. In Proceedings of ACM Symposium on Principles of Programming Languages, pages 119–130, 1978.

[12] J.R. Hindley and J.P. Seldin. Introduction to Combinators and λ-calculus. London Mathematical Society, Student Texts. Cambridge University Press, 1986.

[13] C.A.R. Hoare. Quicksort. Computer Journal, 5(1), 1962.

136 BIBLIOGRAPHY

[14] W.A. Howard. The formulae-as-type notion of construction, pages 479–490. Academic Press, 1980.

[15] P. Hudak and P. Wadler. Report on the programming language Haskell. Technical Report YALEU/DCS/RR-777, Yale University, 1990.

[16] T. Johnsson. Efficient compilation of lazy evaluation. In Proceedings of ACM Conference on Compiler Construction, pages 58–69, 1984.

[17] J.-L. Krivine. Lambda-calcul, types et modèles. Etudes et recherches en informatique. Masson, 1990.

[18] P. Landin. The next 700 programming languages. In Communications of the ACM, volume 9, pages 157–164, 1966.

[19] X. Leroy. The ZINC experiment: an economical implementation of the ML language. Technical Report 117, INRIA, 1990.

[20] X. Leroy. The Caml Light system, release 0.6 — Documentation and user’s manual. Technical report, INRIA, 1993. Included in the Caml Light 0.6 distribution.

[21] X. Leroy and P. Weis. Manuel de référence du langage Caml. InterÉditions, 1993.

[22] J. MacCarthy. Lisp 1.5 Programmer’s Manual. MIT Press, Cambridge, Mass., 1962.

[23] D. MacQueen. Modules for Standard ML. In Proceedings of the ACM Conference on Lisp and Functional Programming, 1984.

[24] M. Mauny. Functional programming using CAML. Technical Report 129, INRIA, 1991.

[25] M. Mauny and D. de Rauglaudre. Parsers in ML. In Proceedings of the ACM Conference on Lisp and Functional Programming, 1992.

[26] M. Mauny and A. Suárez. Implementing functional languages in the categorical abstract machine. In Proceedings of the ACM International Conference on Lisp and Functional Programming, pages 266–278, 1986.

[27] R. Milner. A theory of type polymorphism in programming. J. Comput. Syst. Sci., 17:348–375, 1978.

[28] R. Milner. A proposal for Standard ML. In Proceedings of the ACM Conference on Lisp and Functional Programming, 1987.

[29] S.L. Peyton-Jones. The Implementation of Functional Programming Languages. Series in Computer Science. Prentice-Hall International, 1987.

[30] S.L. Peyton-Jones. Mise en œuvre des langages fonctionnels de programmation. Manuels Informatiques Masson. Masson, 1990.

[31] J. Rees and W. Clinger. The revised^{3} report on the algorithmic language Scheme. In SIGPLAN Notices, volume 21, 1987.

BIBLIOGRAPHY 137

[32] J. A. Robinson. Computational logic: the unification computation. In Machine Intelligence, volume 6 of American Elsevier. B. Meltzer and D. Mitchie (Eds), 1971.

[33] R. Sedgewick. Algorithms. Addison Wesley, 1988.

[34] D. Turner. SASL language manual. Technical report, St Andrews University, 1976.

[35] D. Turner. Recursion equations as a programming language, pages 1–28. Cambridge University Press, 1982.

[36] D. Turner. Miranda: a non-strict functional language with polymorphic types. In Functional Programming Languages and Computer Architecture, number 201 in Lecture Notes in Computer Science, pages 1–16. Springer Verlag, 1985.

[37] P. Weis, M.V. Aponte, A. Laville, M. Mauny, and A. Suárez. The CAML reference manual. Technical Report 121, INRIA, 1990.

[38] P. Weis and X. Leroy. Le langage Caml. InterÉditions, 1993.